If I is the current in the external resistor R and r is the internal resistance of the cell, the equation for the whole circuit is:

            E = I(R + r)          …….(i)

            For the same current in the external resistor R, the terminal voltage is given by:

            V = IR                  …….(ii)

Use equations (i) and (ii) to solve for r by eliminating I and write this equation the form:

y = mx + b

            Identify the independent variable and the dependent variable.   

   Draw an appropriate graph and measure its slope.   

Use the slope of the graph to obtain a value of the internal resistance r of the cell.

   i need helpppppppppppppp how the graph looks like ????   
Answer #1

r = PD/(V/R) PD =[ E -- V]
r = 2.22/(4,43/1)
r =2.22/4,43
r= 1,5 ohms approx

V is the terminal voltage of supply due to current flow through external resistance
therefore the potential difference of EMF and voltage is due to current flow through
the internal resistance (r)